### v1

# FlowDubber: Movie Dubbing with LLM-based Semantic-aware Learning and Flow Matching based Voice Enhancing

Abstract.
Movie Dubbing aims to convert scripts into speeches that align with the given movie clip in both temporal and emotional aspects while preserving the vocal timbre of a given brief reference audio. Existing methods focus primarily on reducing the word error rate while ignoring the importance of lip-sync and acoustic quality. To address these issues, we propose a large language model (LLM) based flow matching architecture for dubbing, named FlowDubber, which achieves high-quality audio-visual sync and pronunciation by incorporating a large speech language model and dual contrastive aligning while achieving better acoustic quality via the proposed voice-enhanced flow matching than previous works. First, we introduce Qwen2.5 as the backbone of LLM to learn the in-context sequence from movie scripts and reference audio. Then, the proposed semantic-aware learning focuses on capturing LLM semantic knowledge at the phoneme level. Next, dual contrastive aligning (DCA) boosts mutual alignment with lip movement, reducing ambiguities where similar phonemes might be confused. Finally, the proposed Flow-based Voice Enhancing (FVE) improves acoustic quality in two aspects, which introduces an LLM-based acoustics flow matching guidance to strengthen clarity and uses affine style prior to enhance identity when recovering noise into mel-spectrograms via gradient vector field prediction. Extensive experiments demonstrate that our method outperforms several state-of-the-art methods on two primary benchmarks. The demos are available at https://galaxycong.github.io/LLM-Flow-Dubber/.
Movie Dubbing, Visual Voice Cloning, Flow Matching
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: arXiv Preprint; April, 2025; arXiv.org
1. Introduction
Movie Dubbing, also known as Visual Voice Cloning (V2C) (Chen et al., 2022), aims to generate a vivid speech from scripts using a specified timbre conditioned by a single short reference audio while ensuring strict audio-visual synchronization with lip movement from silent video, as shown in Figure 1. It attracts great attention in the multimedia community and promises significant potential in real-world applications such as film post-production and personal speech AIGC.
Figure 1. (a) Illustration of V2C task. (b) Illustration of the proposed FlowDubber. It brings a new level of high-quality lip-sync and pronunciation by incorporating a large speech language model and dual contrastive aligning while achieving better acoustic quality via voice-enhanced flow matching.
Previous dubbing works (Zhang et al., 2024, 2025a; Cong et al., 2024b, 2023; Chen et al., 2022) achieve significant progress in improving pronunciation and are dedicated to reducing the word error rate (WER) of generated speech. They can be mainly divided into two groups. Since the dubbing resources are limited in scale (copyright issues) and are always accompanied by background sounds or environmental noise, one class of methods (Zhang et al., 2024, 2025a) focuses primarily on leveraging external knowledge to improve pronunciation clarity by pre-training on clear large-scale text-to-speech corpus (Zen et al., 2019). For example, Speaker2Dubber (Zhang et al., 2024) proposes a two-stage dubbing architecture, which allows the model to first learn pronunciation via multi-task speaker pre-training on Libri-TTS 100 dataset and then optimize duration in stage two. Then, by pre-training on larger TTS corpus Libri-TTS 460 dataset, ProDubber (Zhang et al., 2025a) proposes another novel two-stage dubbing method based on Style-TTS2 model (Li et al., 2023), including prosody-enhanced pre-training and acoustic-disentangled prosody adapting. However, these pre-training methods rely too much on the TTS architecture (Ren et al., 2021; Li et al., 2023) and mainly adopt a Duration Predictor (DP) (Cong et al., 2024b) to produce rough duration without considering intrinsic relevance with lip motion, resulting in poor audio-visual sync.
The other family of methods (Chen et al., 2022; Cong et al., 2023, 2024b) do not care about pre-training, but try to decline WER by associating other related modality information that helps with pronunciation. For example, Styledubber (Cong et al., 2024b) proposes a multi-modal style adaptor to learn pronunciation style from the reference audio and generate intermediate representations informed by the facial emotion presented in the video. However, due to the introduction of time stretching, StyleDubber (Cong et al., 2024b) can only keep the global time alignment (i.e., the total length of the synthesized dubbing is consistent with the target), which is still unsatisfactory in fine-grained matching with lip motion, bringing a bad audio-visual experience.
Except for the alignment issues mentioned above, the existing dubbing methods suffer from acoustic quality degradation, even in the advanced two-stage dubbing pre-training methods. For example, Speaker2Dubber (Zhang et al., 2024) freezes the text encoder in the second stage, which helps to maintain pronunciation. However, its use of a traditional FastSpeech2-based (Ren et al., 2021) transformer fails to handle the complex and diverse spectrum changes, leading to subpar acoustic quality. In addition, the authoritative acoustic quality measurement predictor UTMOS (Saeki et al., 2022) reveals that the acoustic quality of current dubbing methods still requires improvement.
Recent advances in speech tokenization (van den Oord et al., 2017; Du et al., 2024b; Mentzer et al., 2024; Ji et al., 2024a) have revolutionized TTS synthesis by bridging the fundamental gap between continuous speech signals and token-based large language models (LLM). Due to LLM demonstrating excellent capability in sequential modeling and contextual understanding, these LLMs-based speech synthesis models achieve human-level expressive and naturalness (Du et al., 2024b; Ye et al., 2025; Wang et al., 2025; Guo et al., 2024b). However, they are struggling to deal with dubbing task. Although some speed-controllable LLM speech models have been proposed, they still lack visual understanding capabilities, and the synthesized speech struggles to align with the lip motion changing in video. Besides, some speech models focus too much on the naturalness of speech, resulting in poor cloning ability, which makes it challenging to maintain speaker similarity in speech synthesis.
To address these issues, we propose an LLM-based flow matching architecture for dubbing, named FlowDubber, which guarantees high-quality audio-visual sync and pronunciation by incorporating a large speech language model and dual contrastive aligning while achieving state-of-the-art acoustic quality via the proposed voice-enhanced flow matching (as shown in Figure 1 (b)). Specifically, we first design an LLM-based Semantic-aware Learning (LLM-SL), which consists of pre-trained textual LLM Qwen2.5-0.5B to model the in-context sequence from movie scripts and reference audio and semantic-aware phoneme learning focuses on capturing the relevance between phoneme pronunciation unit and LLM semantic knowledge. Then, the proposed dual contrastive aligning (DCA) ensures mutual alignment between lip movement and phoneme sequence, reducing ambiguities where similar phonemes might be confused. Finally, we propose a novel Flow-based Voice Enhancing (FVE) module, which improves the acoustic quality from two sub-components: LLM-based acoustics flow matching guidance and style flow matching prediction. The LLM-based acoustics flow matching guidance focuses on improving the clarity of acoustics during recovering noise to mel-spectrograms by gradient vector field prediction conditioned on LLM.
The main contributions of the paper are as follows:
  • We propose a powerful dubbing architecture FlowDubber, which incorporates LLM for semantic learning and flow matching for acoustic modeling to enable high-quality dubbing, including lip-sync, acoustic clarity, speaker similarity.
  • We devise an LLM-based Semantic-aware Learning (LLM-SL) to absorb token-level semantic knowledge, which is convenient to achieve precisely lip-sync for dubbing by associating proposed dual contrastive aligning.
  • We design a Flow-based Voice Enhancing mechanism to enhance the semantic information from LLM, refining the flow-matching generation process for high speech clarity.
  • Extensive experimental results demonstrate the proposed FlowDubber  performs favorably against state-of-the-art models on two dubbing benchmark datasets.
2. Related Work
2.1. Visual Voice Cloning
The V2C task (Chen et al., 2022) requires generating high-quality dubbing speech how a text might be said, but in step with the lip movements portrayed by video character, and in vocal style exemplified by reference audio (Kim et al., 2025; Sung-Bin et al., 2025; Choi et al., 2025, 2024; Ye and Shan, 2025; Zhang et al., 2025b; Zheng et al., 2025; Li et al., 2025). Some works focus primarily on improving the pronunciation clarity (Cong et al., 2023, 2024b; Zhao et al., 2024). For example, SOTA dubbing method ProDubber (Zhang et al., 2025a) and Speaker2Dubber (Zhang et al., 2024) propose a pre-training framework to learn clear pronunciation representation from a large-scale text-to-speech corpus (Zen et al., 2019). However, they over-rely on the TTS architecture and use an inaccurate duration predictor (Zhang et al., 2025a) to estimate the lip speaking time, without considering the intrinsic connection between visual movement and speech content. Besides, StyleDubber (Cong et al., 2024b) attempts to use time stretching to balance articulation and lip-sync. Although the overall length can remain consistent, this does not fundamentally achieve the audio-visual alignment in fine-grained lip-sync. In this work, we propose FlowDubber, a novel dubbing architecture that combines LLM-based semantic-aware learning with dual contrastive alignment to achieve high-quality lip synchronization, and the proposed flow-matching enhancing mechanism delivers better acoustic quality than existing dubbing methods.
Figure 2. Overall framework of FlowDubber. It consists of LLM-based Semantic-aware Learning (LLM-SL), lip-phoneme Dual Contrastive Aligning (DCA), and Flow-based Voice Enhancing (FVE). Specifically, the LLM-SL includes Qwen2.5-0.5B speech language model and semantic-aware phoneme learning to keep pronunciation while aligning with DCA. The FVE consists of style flow matching prediction and LLM-based acoustics flow matching guidance to improve the acoustics quality.
2.2. Large Language Model and Speech Codec
The remarkable success of Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2025; Yang et al., 2024b) and the autoregressive (AR) model brings significant advancements in the field of speech synthesis. VALL-E (Wang et al., 2023a) first converts speech into neural codec tokens and treats the speech synthesis as a next-token prediction task. Subsequently, extensive research focuses on speech codecs and LLM-based speech generators to improve the synthesis performance. For example, DAC (Kumar et al., 2023) adopts the residual vector quantization and the multi-scale STFT discriminators to obtain higher-quality discrete speech tokens. Wavtokenizer (Ji et al., 2024b) and X-codec (Ye et al., 2024) further improved the efficiency of codec and addressed the semantic shortcomings of previous codes. Besides, LLM-based speech synthesis systems combine the AR model with other components (Anastassiou et al., 2024; Chen et al., 2024a) or rely on continuous acoustic features (Meng et al., 2024; Zhu et al., 2024) to achieve better performance. Recently, Llasa (Ye et al., 2025) investigated the effects of training-time inference-time scaling in LLM-based speech synthesis. However, they still lack visual understanding capability, and the generated speech struggles to align with the lip movement. In this paper, we propose powerful dubbing model that can achieve the best lip-sync and inherit the acoustic knowledge of LLM via flow-matching learning and phoneme learning, achieving advanced results against all dubbing methods.
2.3. Speech Synthesis and Flow Matching
Flow Matching (Lipman et al., 2022) is a simulation-free approach to training continuous normalizing flow (Chen et al., 2018) models, capable of modeling arbitrary probability paths and capturing the trajectories represented by diffusion processes (Song et al., 2021). Due to the high quality and faster speed, flow matching has attracted significant attention in speech generation (Guo et al., 2024a; Kim et al., 2023; Le et al., 2023). For example, Matcha-TTS (Mehta et al., 2024) adopts the optimal transport conditional flow matching in single speaker TTS synthesis and Stable-VC (Yao et al., 2024) adopts it in voice conversion field to improve fidelity. F5-TTS (Chen et al., 2024b) is another powerful TTS model to reconstruct high-quality mel-spectrograms by flow matching. Then, CosyVoice 2.0 (Du et al., 2024a, b) has further proven its superior performance by combining flow matching with LLM. However, these methods are not suited to V2C dubbing task due to they inability to perceive proper pause in step with lip motion. Recently, EmoDub (Cong et al., 2024a) introduces classifier guidance in flow matching to control emotions via input labels and intensity. In contrast, after integrating semantic-aware phoneme learning and lip-motion aligning, we focus on refining the flow-matching generation process to ensure clarity by introducing semantic knowledge from LLM via classifier-free guidance.
3. Methods
3.1. Overview
The target of the overall movie dubbing task is:
(1) $\hat{{Y}}=\mathrm{FlowDubber}(W_{r},T_{c},V_{s}),$
where the $V_{s}$ represents the given a silent video clip, $W_{r}$ is a reference waveform used for voice cloning, and $T_{c}$ is current piece of text to convey speech content. The goal of FlowDubber  is to generate a piece of high-quality speech $\hat{Y}$ that guarantees precise lip-sync with silent video, high speaker similarity, and clear pronunciation. The main architecture of the proposed model is shown in Figure 2. Specifically, we introduce pre-trained textual LLM Qwen2.5-0.5B as the backbone of the speech language model to model the in-context sequence from movie scripts and reference audio by discretizing them. Then, the semantic knowledge of speech tokens is adapted to the phoneme level by semantic-aware phoneme learning. Next, the proposed Dual Contrastive Aligning (DCA) ensures the cross model alignment between lip-motion and phoneme level information from LLM. Finally, Flow-based Voice Enhancement (FVE) enhances the fused information from two aspects: Style Flow Matching Prediction aims to keep the speaker similarity and LLM-based Acoustics Flow Matching Guidance focuses on improving the acoustics clarity and suppressing noise. We detail each module below.
3.2. LLM-based Semantic-aware Learning
Different from the previous dubbing works (Cong et al., 2023; Zhao et al., 2024), we introduce LLM-based semantic-aware learning to capture the phoneme level pronunciation via the powerful in-context learning capabilities of LLM (Qwen2.5-0.5B) between text token in movie script and semantic and identity token in reference audio.
Speech Tokenization. This module aims to transform the speech signal of reference audio $R_{a}$ into a sequence of semantic tokens ${h}_{q}$. It first utilizes a pre-trained self-supervised learning (SSL) model, wav2vec 2.0 (Baevski et al., 2020), to translate speech signals into a semantic embedding sequence. Then, the semantic encoder $S_{encoder}(\cdot)$, constructed with 12 ConvNeXt (Liu et al., 2022) blocks and 2 downsampling blocks, is employed to process and down-sample the sequence further into an encoding sequence $h$:
(2) ${H}_{q}=\text{VQ}({h}),{h}=S_{encoder}(\text{wav2vec2.0}(R_{a})),$
where the output ${H}_{q}$ represents semantic tokens from $h$ by Vector Quantization $\text{VQ}(\cdot)$ layers. Specifically, the $\text{VQ}(\cdot)$ adopts factorized codes manner and has an 8192 codebook size and 8 codebook dimensions. Different from CosyVocie 2.0 (Du et al., 2024b), we also extract global tokens ${G}_{q}$ from reference audio’s mel-spectrogram by Finite Scalar Quantization (FSQ) (Mentzer et al., 2024) to keep the speaker characteristics (Wang et al., 2025).
Speech Language Model. Inspired by LLM successes, we employ the pre-trained textual LLM Qwen2.5-0.5B (Yang et al., 2024b) as the backbone of the speech language model. Specifically, we formulate GPT (Radford et al., 2019) architecture as the next-token prediction paradigm, which adopts a decoder-only autoregressive transformer architecture:
(3) $P(o_{1:N_{o}})=\prod_{i=1}^{N_{o}}P(o_{i}|T_{q},H_{q},G_{q},o_{1},\cdots,o_{i-1}),$
where $o_{i}$ is the i-th generated speech token, and $N_{o}$ is the length of generated speech tokens. The $T_{q}$ represents text tokens by converting raw text $T_{c}$ using a byte pair encoding (BPE)-based tokenizer. $H_{q}$ are semantic tokens and $G_{q}$ are global tokens from reference audio. By inputting the concatenation of $T_{q}$, $G_{q}$, $H_{q}$ and previous special tokens $(o_{1},...,o_{i-1})$, model can autoregressively generate current speech tokens $o_{i}$ with in-context semantic knowledge.
Phoneme Level Semantic-aware Module. Compared with zero-shot TTS, movie dubbing must be strictly matched with lip movements from silent video to achieve audio-visual synchronization. The proposed phoneme-level semantic-aware module aims to capture the semantic knowledge from the speech language model at the phoneme level, which helps preserve pronunciation and enables fine-grained alignment between phoneme unit and lip motion sequence. Specifically, the phoneme-level semantic-aware module consists of cross-modal transformers $\hat{Z}^{[i]}_{S\rightarrow P}$ to calculate the relevance between textual phoneme embedding and LLM speech knowledge, which can be formulate as:
(4) $\begin{split}\hat{Z}^{[i]}_{S\rightarrow P}=&\;\text{LLM}^{[i],\text{mul}}_{S\rightarrow P}(\text{LN}(Z^{[i-1]}_{S\rightarrow P}),\text{LN}(Z^{[0]}_{S}))+\text{LN}(Z^{[i-1]}_{S\rightarrow P}),\\ Z^{[i]}_{S\rightarrow P}=&\;f_{\theta^{[i]}_{S\rightarrow P}}(\text{LN}(\hat{Z}^{[i]}_{S\rightarrow P})+\text{LN}(\hat{Z}^{[i]}_{S\rightarrow P}),\end{split}$
where ${\text{LN}(\cdot)}$ denotes the layer normalization in cross modal transformer, $i=\{1,...,D\}$ denotes the number of feed-forwardly layers, and $f_{\theta}$ is a position-wise feed-forward sublayer parametrized by $\theta$. ${\text{LLM}^{[i],\text{mul}}_{A\rightarrow L}}(\cdot)$ is a multi-head attention as follows:
(5) $\text{LLM}^{[i],\text{mul}}_{S\rightarrow P}=\mathrm{softmax}(\frac{{E_{pho}(\text{G2P}(T_{c}))}{S_{llm}}^{\top}}{\sqrt{d_{m}}}){S_{llm}},$
[…]
3.3. Dual Contrastive Aligning for Dubbing
This module is designed to solve alignment problems in dubbing by introducing a Dual Contrastive Learning (DAL) between lip movement sequence and phoneme sequence.
Lip-motion Feature Extractor. To ensure fairness for measuring the alignment ability of DAL, we first use the same extractor (Cong et al., 2023; Zhang et al., 2024, 2025a) to obtain lip motion features from silent videos $V_{s}$:
(6) $z_{m}=\mathrm{LipEncoder}(\mathrm{LipCrop}(V_{s})),$
where $z_{m}\in\mathbb{R}^{{L_{v}}\times{d_{m}}}$ denotes the output lip motion embedding, $L_{v}$ indicates the length of lip sequence, and $d_{m}$ is embedding size. The $\mathrm{LipCrop}(\cdot)$ uses the face landmarks tool to crop mouth area and $\mathrm{LipExtra}(\cdot)$ consists of 3D convolution, ResNet-18, and 1D convolution (Cong et al., 2023) to capture dynamic lip-motion representation.
Dual Contrastive Learning. We focus on learning the intrinsic correlation between phoneme-level pronunciation and lip movement to achieve reasonable alignment for movie dubbing. Following the contrastive learning manner, we introduce the InfoNCE loss (Tian, 2022) to encourage the model to distinguish correct lip-phoneme pairs. Specifically, we first treat the lip motion features $z_{m}$ as queries and the phoneme embeddings $z_{p}$ as keys. To establish positive pairs, we align each lip motion frame with its corresponding phoneme based on ground-truth timing annotations by MFA and FPS. This ensures that each $z_{m}^{i}$ should be maximally similar to its temporally aligned $z_{p}^{j}$, while being distinct from other phonemes:
(7) $\mathcal{L}_{mp}=-\sum_{i}\log\frac{\sum_{j\in+}\exp(z_{m}^{i}\cdot z_{p}^{j}/\tau)}{\sum_{j}\exp(z_{m}^{i}\cdot z_{p}^{j}/\tau)},$
where $i\in[0,L_{v}-1]$ represents the $i$-th frame of the lip sequence and $j\in[0,L_{t}-1]$ represents the $j$-th textual phoneme from whole sequence. The + means positive sample pairs, which are calculated in advance based on the ground-truth information during training (Cong et al., 2024a). Conversely, we introduce a second contrastive loss by reversing the roles: treating phoneme features $z_{p}$ as queries and lip motion embeddings $z_{l}$ as keys. In this case, each phoneme seeks to retrieve its temporally aligned lip feature while suppressing mismatched lip frames:
(8) $\mathcal{L}_{pm}=-\sum_{j}\log\frac{\sum_{i\in+}\exp(z_{p}^{j}\cdot z_{m}^{i}/\tau)}{\sum_{i}\exp(z_{p}^{j}\cdot z_{m}^{i}/\tau)},$
Unlike single-directional contrastive learning, which only aligns one modality to the other, the proposed DCA ensures mutual alignment, reducing ambiguities where similar phonemes might be confused. Finally, we simply use their average as dual contrastive loss:
(9) $\mathcal{L}_{dua}=\frac{1}{2}\mathcal{L}_{mp}+\frac{1}{2}\mathcal{L}_{pm}.$
Aligning Phoneme Level Feature. The similarity matrix between textual phoneme embedding and lip movement embedding ${Sim}(z_{m},z_{p})$ is constrained by dual contrastive learning, then the ${Sim}(z_{m},z_{p})$ further guidance to the generation of aligning sequences, including: (1) lip-related aligning sequences $C_{lip}$. (2) phoneme-related aligning sequences. Specifically, the $C_{lip}$ is obtained by multi-head attention module in  (Cong et al., 2023), in which the $z_{p}$ serves as key and value, and the $z_{m}$ is the query. Unlike (Cong et al., 2023), the learnable ${Sim}(z_{m},z_{p})$ is used as multi-head attention weight matrix to provide correct relevance. Next, by monotonic alignment search (MAS) (Kim et al., 2020), the ${Sim}(z_{m},z_{p})\in\mathbb{R}^{{L_{v}}\times{L_{t}}}$ is flat to mapping table $tab\in\mathbb{R}^{{L_{t}}\times{1}}$, which records the number of video frames corresponding to each phoneme unit. Finally, the $tab$, $LLM_{p}$, $z_{p}$, and $C_{lip}$ are associated to mel-spectrograms level prior conditions $\mu$:
(10) $\mu=\mathcal{F}([C_{lip},\mathrm{Up}(LLM_{p},z_{p},tab)),$
where $\mathrm{Up}(\cdot)$ is used to expand $LLM_{p}$ and $z_{p}$ to video level according to mapping $tab$. The $\mathcal{F}(\cdot)$ indicates the fusion module, which consists of two 2D upsampling convolutional layers and transformer-based mel-decoder. The output $\mu\in\mathbb{R}^{{L_{m}}\times{d_{m}}}$, where ${l_{m}}$ and $d_{m}$ represent the length and embedding size of mel-spectrogram.
3.4. Flow-based Voice Enhancing
In this section, we introduce flow-based voice enhancing, including Style Flow Matching Prediction to inject speaker style into flow matching and LLM-based Acoustics Flow Matching Guidance to improve the clarity of generated speech by enhancing semantic information from the LLM.
Style Flow Matching Prediction. Flow matching generates mel-spectrograms $\hat{M}$ from Gaussian noise by a vector field. Given mel-spectrogram space with data $M$, where $M\sim q(M)$. We aim to train a flow matching network to fit $q(M)$ by predicting the probability density path given the vector field, which can be defined as $p_{t}(x)$. Here $t\in[0,1]$, $p_{0}(x)=\mathcal{N}(x;0,I)$ and $p_{1}(x)=q(x)$. Flow matching can predict the probability density path, gradually transforming $x_{0}\sim p_{0}(x)$ into $M\sim q(M)$. Our Flow matching prediction network is based on optimal-transport conditional flow matching (OT-CFM). OT-CFM uses a linear interpolation flow $\phi_{t}(x)=(1-(1-\sigma_{\min})t)x_{0}+tM$, which satisfies the marginal condition $\phi_{0}(x)=x_{0}$ and $\phi_{1}(x)=M$. The gradient field vector field of OT-CFM is $u_{t}(\phi_{t}(x)|M)=M-(1-\sigma_{min})x_{0}$. The training objective of Flow matching prediction network is to predict the gradient vector field $v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)$, which should be close to $u_{t}(\phi_{t}(x)|M)$: Here $\mu_{SATL}$ is style-enhanced mel-spectrograms level prior according to $\mu$ in Eq. 10. To enhance speakers’ style, we introduced SATL in flow matching. Specifically, during the flow matching generation process, SATL introduces and enhances style information through affine transformation, which can be formulated as:
(11) $\mu_{SATL}=\gamma_{2}(\gamma_{1}\mu+\beta_{1})+\beta_{2},$
where $\gamma_{1},\gamma_{2},\beta_{1},\beta_{2}$ are parameters predicted by SATL based on style features. We train Style Flow Matching Prediction Network using condition $\mu_{SATL}$. We aim for the Flow Matching prediction network to generate the target mel-spectrogram M conditioned on a given $\mu_{SATL}$. During the inference process, Flow Matching prediction network solves the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)dt$ from $t=0$ to $t=1$ to generate a mel-spectrogram $\hat{M}$.
LLM-based Acoustics Flow Matching Guidance. To enhance the clarity of the generated audios, we enhanced the mel-spectrograms level prior conditions by LLM-based Acoustics Flow Matching Guidance. We observed that the generation process in LLM includes semantic tokens and text tokens, which introduce semantic knowledge. Specifically, we enhance LLM’s information in flow matching process to improve speech clarity based on classifier-free guidance, which can be formulated as:
$\tilde{v}_{t}(\phi_{t}(x)|\mu,\theta)=v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)+\alpha\Big{(}v_{t}(\phi_{t}(x)|\mu_{SATL},\theta)-v_{t}(\phi_{t}(x)|\mu^{\prime},\theta)\Big{)}$
Here $\mu^{\prime}=\mathcal{F}([C_{lip},\mathrm{Up}(\phi,z_{p},tab))$, $\phi$ refers to zero vector. By enhancing only the LLM information to improve speech clarity with classifier-free guidance, we can control the mel-spectrograms clarity by removing noise and boosting overall quality. As a result, the proposed guidance mechanism strengthens the semantic information accessible to the flow-matching prediction network, thereby refining the gradient vector field generation process to achieve higher speech clarity.
Table 1. Compared with related Dubbing methods on Chem benchmark. For the Dub 1.0 setting, we use the ground truth audio as reference audio, for the Dub 2.0 setting, we use the non-ground truth audio from the same speaker within the dataset as the reference audio which is more aligned with practical usage in dubbing. M2CI-Dub (Zhao et al., 2025) has no provide setting2 audio.
Setting Dubbing Setting 1.0 Dubbing Setting 2.0
Methods LSE-C $\uparrow$ LSE-D $\downarrow$ SIM-O $\uparrow$ WER $\downarrow$ UTMOS $\uparrow$ LSE-C $\uparrow$ LSE-D $\downarrow$ SIM-O $\uparrow$ WER $\downarrow$ UTMOS $\uparrow$
GT 8.12 6.59 0.927 3.85 4.18 8.12 6.59 0.927 3.85 4.18
Imagin (Lee et al., 2023a) (ICASSP 2023) 1.98 12.50 0.250 62.24 2.62 1.96 12.53 0.184 68.13 2.13
V2C-Net (Chen et al., 2022) (CVPR 2022) 1.97 12.17 0.154 90.47 1.81 1.82 12.09 0.087 94.59 1.76
HPMDubbing (Cong et al., 2023) (CVPR 2023) 7.85 7.19 0.536 16.05 2.16 3.98 9.50 0.187 29.82 2.01
StyleDubber (Cong et al., 2024b) (ACL 2024) 3.87 10.92 0.607 13.14 3.02 3.74 11.00 0.540 14.18 3.04
Speaker2Dubber (Zhang et al., 2024) (MM 2024) 3.76 10.56 0.663 16.98 3.61 3.45 11.17 0.583 18.10 3.64
M2CI-Dub (Zhao et al., 2025) (ICASSP 2025) 7.99 6.91 0.621 12.85 3.15 - - - - -
EmoDubber (Cong et al., 2024a) (CVPR 2025) 8.11 6.92 0.718 11.72 3.82 8.09 6.96 0.625 12.81 3.75
Produbber (Zhang et al., 2025a) (CVPR 2025) 2.58 12.54 0.387 9.45 3.85 2.78 12.14 0.310 11.69 3.76
Ours ($\alpha$ = 0.0) 8.21 6.89 0.754 9.96 3.91 8.17 6.96 0.648 12.95 3.89
Table 2. The zero shot results under Dub 3.0 setting, which use unseen speaker as refernce audio.
Methods LSE-C $\uparrow$ LSE-D $\downarrow$ WER $\downarrow$ UTMOS $\uparrow$
StyleDubber (Cong et al., 2024b) 6.17 9.11 15.10 3.50
Speaker2Dubber (Zhang et al., 2024) 4.83 10.39 15.91 3.53
ProDubber (Zhang et al., 2025a) 5.49 9.49 14.25 3.94
Ours ($\alpha$ = 0.0) 7.43 6.64 13.96 3.98
4. Experimental Results
4.1. Implementation Details
The semantic tokenizer consists of 12 ConvNeXt blocks and 2 downsampling blocks. The codebook size of VQ is 8192. The ECAPA-TDNN in the global tokenizer features an embedding dimension of 512, while the GE2E in style flow matching prediction is used to extract speaker embedding with dimension of 256. We follow the Qwen2.5 tokenizer to process raw text. The cross model transformer is consists of 8 layers with 2 heads, and the dimension size is 256. The solver of conditional flow matching is euler, and we set the number of ODE steps is 10. In dual contrastive aligning, we use 4 heads for multi-head attention with 256 hidden sizes to obtain the attention similarity matrix. The temperature coefficient $\tau$ of $\mathcal{L}_{pm}$ and $\mathcal{L}_{mp}$ as both 0.1. We remove the vocoder bias denoiser in HiFiGAN. In data process, the video frames are sampled at 25 FPS and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following (Martinez et al., 2020; Ma et al., 2021). The window length, frame size, and hop length in STFT are 640, 1,024, and 160, respectively. For LLM-based Voice Enhancement Guidance, the guidance scale is set between 0.0 and 0.8 empirically. We set the batch size to 16 on Chem dataset and 64 on GRID. Our model is implemented in PyTorch. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU.
4.2. Datasets
We choose a real-person dubbing dataset to conduct extensive experiments to reasonably evaluate lip-sync. Our dataset mainly includes Chem and GRID.
Chem is a popular dubbing dataset recording a chemistry teacher speaking in the class (Prajwal et al., 2020a). It is collected from YouTube, with a total video length of approximately nine hours. For complete dubbing, each video has clip to sentence-level (Hu et al., 2021). The number of train, validation, and test data are 6,082, 50, and 196, respectively.
GRID is a dubbing benchmark for multi-speaker dubbing (Cooke et al., 2006). The whole dataset has 33 speakers, each with 1,000 short English samples. All participants are recorded in studio with unified background. The number of train and test data are 32,670 and 3,280.
4.3. Evaluation Metrics
We adopt multiple metrics to comprehensively evaluate the acoustic quality of synthesized dubbing. We abandon some old evaluation metrics and follow the latest speech synthesis technology to evaluate the synthesis quality. We use LSE-C/D instead of MCD-DTW-SL to evaluate lip-sync. We use SIM-O instead of SECS to evaluate speaker similarity. We adopt UTMOS instead of MCD-DTW to evaluate quality of speech. Below is the details of each metrics:
LSE-C and LSE-D. To evaluate the synchronization between the generated speech and the video quantitatively, we adopt Lip Sync Error Distance (LSE-D) and Lip Sync Error Confidence (LSE-C) as our metrics, which are widely used to lip reading (Yemini et al., 2024), talking face (Wang et al., 2023b; Jang et al., 2024), and video dubbing task (Hu et al., 2021; Lu et al., 2022). These metrics are based on the pre-trained SyncNet (Chung and Zisserman, 2016), which can explicitly test for lip synchronization in unconstrained videos in the wild (Prajwal et al., 2020b; Chung and Zisserman, 2016). Compared to the length metric MCD-SL (Chen et al., 2022), we believe that LSE-C and LSE-D can more accurately measure the synchronization of vision and audio. The discussion of the two kinds of metrics is in Appendix.
SIM-O. To evaluate the timbre consistency between the generated dubbing and the reference audio, we employ the SIM-O following (Ju et al., 2024) to compute the similarity of speaker identity. The similarity score predicted by WavLMTDNN is in the range of [-1; 1], where a larger value indicates a higher similarity of input samples.
UTMOS. UTMOS (Saeki et al., 2022) is an authoritative acoustic predictor, which focuses on evaluating the acoustic quality of synthesized speech (Zhang et al., 2025a; Ju et al., 2024; Wang et al., 2024; Du et al., 2024b; Ye et al., 2025; Wang et al., 2025), particularly by assessing naturalness, intelligibility, prosody, and expressiveness.
DNSMOS. Deep Noise Suppression MOS (DNSMOS) (Reddy et al., 2021) is designed to assess the quality of speech processed by noise suppression algorithms, measuring clarity, naturalness, background noise quality, and overall quality.
SNR score. The signal-to-noise ratio (SNR) score is a deep learning-based estimation system (Li et al., 2020) to assess the clarity of speech. A larger SNR corresponds to higher speech clarity.
WER. The Word Error Rate (WER) (Morris et al., 2004) is used to measure pronunciation accuracy by using Whisper-V3 (Radford et al., 2023) as the ASR model.
Table 3. Compared with related Dubbing methods on GRID benchmark under the same dub setting as the Chem benchmark.
Setting Dubbing Setting 1.0 Dubbing Setting 2.0
Methods LSE-C $\uparrow$ LSE-D $\downarrow$ SIM-O $\uparrow$ WER $\downarrow$ UTMOS $\uparrow$ LSE-C $\uparrow$ LSE-D $\downarrow$ SIM-O $\uparrow$ WER $\downarrow$ UTMOS $\uparrow$
GT 7.13 6.78 0.866 0.00 3.94 7.13 6.78 0.866 0.00 3.94
Imagin (Lee et al., 2023a) (ICASSP 2023) 4.69 10.14 0.424 44.37 2.53 4.55 10.27 0.458 39.15 2.48
V2C-Net (Chen et al., 2022) (CVPR 2022) 5.59 9.52 0.430 47.82 2.41 5.34 9.76 0.356 49.09 2.40
HPMDubbing (Cong et al., 2023) (CVPR 2023) 5.76 9.13 0.461 45.51 2.14 5.82 9.10 0.359 44.15 2.11
StyleDubber (Cong et al., 2024b) (ACL 2024) 6.12 9.03 0.754 18.88 3.73 6.09 9.08 0.617 19.58 3.71
Speaker2Dubber (Zhang et al., 2024) (MM 2024) 5.27 9.84 0.734 17.04 3.69 5.19 9.93 0.606 17.00 3.73
EmoDubber (Cong et al., 2024a) (CVPR 2025) 7.12 6.82 0.802 18.53 3.83 7.10 6.89 0.665 19.75 3.81
Produbber (Zhang et al., 2025a) (CVPR 2025) 5.23 9.59 0.791 18.60 3.87 5.56 9.37 0.663 19.17 3.86
Ours ($\alpha$ = 0.0) 7.27 6.72 0.811 18.54 3.97 7.20 6.75 0.679 19.24 3.95
Table 4. The Clarity performance of using different scale $\alpha$ in acoustics flow matching guidance. Note that DNSMOS, SNR Score, and UTMOS are not subjective metrics.
Guidance Scale DNSMOS $\uparrow$ SNR Score $\uparrow$ UTMOS $\uparrow$
Produbber (Zhang et al., 2025a) 3.664 23.703 3.849
Ours ($\alpha$ = 0.0) 3.745 26.341 3.912
Ours ($\alpha$ = 0.2) 3.777 26.657 3.929
Ours ($\alpha$ = 0.4) 3.799 26.706 3.940
Ours ($\alpha$ = 0.6) 3.819 26.903 3.953
Ours ($\alpha$ = 0.8) 3.829 27.016 3.960
4.4. Comparison with SOTA Dubbing Methods
To prove the effectiveness of the proposed model, we compare the proposed method with the most advanced dubbing models. All experimental results use the official code or providing checkpoint.
Results on the Chem Dataset. As shown in Table 1, our method achieves the best performance on almost all metrics on the Chem benchmark, whether in setting or setting 2. First, our method achieves the best LSE-C and LSE-D, with absolute improvements of 5.63% and 5.65% than the SOTA dubbing method ProDubber (Zhang et al., 2025a), demonstrating the effectiveness of our methods in lip-sync by LLM-based semantic-aware learning and dual contrastive aligning. Furthermore, in the speaker similarity (see SIM-O), our method improves 13.72% on dub setting1 and 11.15% on dub setting2 than Speaker2Dubber (Zhang et al., 2025a). Please note that setting 2 is more challenging than setting 1. Besides, the dubbing synthesis quality of our method is the highest among all dubbing methods, with a UTMOS score of 3.91. In summary, FlowDubber  is a comprehensive dubbing model that makes up for the shortcomings of previous methods in audio-visual synchronization, speaker similarity, and dubbing synthesis quality, and achieves a WER comparable to SOTA.
Results on the GRID Dataset. As shown in Table 3, a similar trend is found in the multi-speaker benchmark. We still achieve SOTA performance in audio-visual synchronization, dubbing synthesis quality, and discrepancy from ground truth in both dubbing settings while maintaining similarity with advanced speaker identity. Specifically, our method can achieve similar WER as ProDubber (Zhang et al., 2025a) while maintaining high audio-visual synchronization: ours LSE-C is absolute improved by 2.03% and LSE-D is absolute improved by 2.87% compared with Produbber, which is extremely important for moving towards automated and high-quality dubbing. Besides, the UTMOS of our method is improved by 12% over Speaker2Dubber (Zhang et al., 2024) on setting 2, which shows that the speech quality synthesized by our dubbing method is the best, even better than the two-stage pre-training manner, which benefits from the proposed voice enhanced flow matching and LLM-based semantic-aware learning.
Results on the Speaker Zero-shot Test. In addition to dubbing benchmarks, we also conduct the zero-shot test to evaluate the generalization performance of models. This setting uses the audio of unseen characters (from another dataset) as reference audio. Here, we use the audio from the Chem dataset as reference audio to measure the GRID dataset. Since there is no target audio at this setting, we only compare LSE-C, LSE-D, WER, and UTMOS for objective evaluation. As shown in Table 2, our proposed method surpasses the current state-of-the-art models and achieves the best performance across all metrics. Specifically, our proposed method achieves the best pronunciation accuracy 13.96% and the best acoustic quality 3.98 than SOTA dubbing method Produbber (Zhang et al., 2025a), even facing the out-of-domain reference audio. Besides, we still achieve the best lip-sync (see LSE-C and LSE-D) in zero-shot setting, which proves the the superiority of the generalization performance of our proposed method.
Table 5. Ablation study of the proposed method on the Chem benchmark dataset with 1.0 setting.
# Methods LSE-C $\uparrow$ LSE-D $\downarrow$ WER $\downarrow$ SIM-O $\uparrow$ UTMOS $\downarrow$
1 w/o FVE 8.18 6.94 13.85 0.620 3.66
2 w/o LLM-SL 8.16 6.95 48.33 0.671 3.76
3 w/o DCA 3.62 10.28 10.04 0.747 3.90
4 w/o Style in FVE 8.19 6.92 14.96 0.582 3.84
5 Full model 8.21 6.89 9.96 0.754 3.91
4.5. Analysis of Voice-Enhanced Flow Matching
To evaluate the effectiveness of LLM-based Acoustics Flow Matching Guidance in improving the clarity of generated speech, we assess the DNSMOS (Reddy et al., 2021), SNR Score and UTMOS metrics across different guidance scale. Table 4 shows the results. As the guidance scale increases, DNSMOS, SNR Score and UTMOS all show improvement, indicating that LLM-based Acoustics Flow Matching Guidance effectively reduces noise and enhances speech clarity, naturalness, and overall quality. Higher scales correlate with better noise suppression, intelligibility, and a more refined listening experience. Besides, DNSMOS increases faster than UTMOS, indicating that LLM-based Acoustics Flow Matching Guidance primarily enhances clarity, as DNSMOS is more closely associated with speech clarity than UTMOS.
Figure 3. The visualization of the mel-spectrograms of ground truth (GT) and synthesized audios obtained by different models. In (a), green arrows point to the video frames that no speak, and green bounding boxes are used to highlight the pauses in the speech. In (b), pink arrows point to the enhanced details of the mel-spectrogram as flow matching guidance scale $\alpha$ increases.
4.6. Ablation Studies
To further investigate the specific effects of main module (LLM-SL, DCA, FVE, Style in FVE) in our proposed method, we conduct ablation studies on the Dub 1.0 setting of the Chem benchmark. The ablation results are presented in Table 5. It shows that all modules contribute significantly to the overall performance, and each module has a different focus. Specifically, when FVE is removed (line 1), UTMOS drops the most, which shows the importance of the proposed voice enhanced flow matching to gradually remove noise and generate high-quality mel-spectrogram. When LLM-SL (line 2)is removed, both WER and UTMOS decrease, with WER being more obvious. This shows that LLM-based semantic-aware learning can provide rich semantic information on phoneme level, which is necessary for clear pronunciation. When removing DCA and using the duration predictor (line 3) to provide alignment, we observe a significant degradation in LSE-C and LSE-D. Although the impact on sound quality is very small (see UTMOS), it is unacceptable for video dubbing. Last, removing Style in FVE has a greater impact on speaker similarity (see SIM-O).
Table 6. Compared with different audio generators. All results build on acoustics flow matching guidance of scale $\alpha$=0.8.
Methods Type LSE-C $\uparrow$ LSE-D $\downarrow$ SIM-O $\uparrow$ UTMOS $\uparrow$
Ours (HiFiGAN) mel. 8.163 6.954 0.745 3.960
Ours (BigVGAN) mel. 8.185 6.932 0.749 3.971
Ours (16K DAC) codec 8.101 6.980 0.703 3.916
Ours (24K CV) codec 8.179 6.958 0.721 4.154
Table 7. Compared with SOTA LLM-based TTS method.
Methods Dub. LSE-C $\uparrow$ LSE-D $\downarrow$ SIM-O $\uparrow$ UTMOS $\uparrow$
CosyVocie 2.0 (Du et al., 2024b) $\times$ 3.001 12.248 0.718 4.252
Llasa-3B (Ye et al., 2025) $\times$ 3.537 11.564 0.662 4.207
Spark-TTS (Wang et al., 2025) $\times$ 2.850 12.347 0.549 4.390
FireRedTTS (Guo et al., 2024b) $\times$ 2.779 12.413 0.529 4.010
Ours (24K CV) $\checkmark$ 8.179 6.958 0.721 4.154
4.7. Compare with Different Audio Generators
Please note that when comparing with the dubbing baseline (Table 1324, and 5), we adopt HiFi-GAN (Kong et al., 2020) as audio generator to convert the mel-spectrogram to waveforms for ensure fairness. Although HiFi-GAN is one of the most popular vocoders, it is not the most advanced. Thus, we explore the upper-bound quality of the generated audio in this section by using different audio generators. Specifically, we select more powerful audio generators: BigVGAN (Lee et al., 2023b), 16K Hz Descript Audio Codec (DAC) (Kumar et al., 2023), and 24K Hz Codec Vocoder (CV) (Du et al., 2024b), respectively. Note that since the DAC and CV do not accept features in the form of mel-spectrogram, we first convert the mel-spectrogram into a waveform by basic HiFi-GAN, then use their codec manner to discretize it, and finally reconstruct the target audio. As shown in Table 6, the results show that 24K CV achieves the best speech quality (see UTMOS), while BigVGAN achieves better alignment and timbre restoration with a slight advantage. Most importantly, we find that all audio generators are better than SOTA dubbing baseline (e.g., Produbber (Zhang et al., 2025a)) or powerful TTS methods (see Table 7) in audio-visual synchronization (see LSE C/D), because the aligning information has been preserved in advance. This is also the advantage of our method, which can be extended by stronger audio generators in the future.
4.8. Compare with LLM-based TTS method
As shown in Table 7, we compare with the recent LLM-based TTS methods. Our method achieves the best performance in LSE-C and LSE-D to maintain synchronization, while ensuring high speech quality. Specifically, our dubbing scheme can approach or even exceed some large-scale TTS methods in UTMOS. For example, our UTMOS is 3.59% higher than FireRedTTS. In contrast, LLM-based TTS methods cannot adapt to dubbing scenes due to the lower LSE-D and LSE-C, proving the bad audio-visual alignment with lip motion. Although Spark-TTS has excellent speech quality, its speaker similarity is poor. In terms of speaker similarity, our method can improve 31.32% on SIM-O than Spark-TTS.
4.9. Qualitative Analysis
We visualize the mel-spectrograms of ground truth and dubbing generated by different models for comparison in Figure 3. The green bounding boxes highlight the pauses in the speech and blue bounding boxes exhibit significant differences in acoustic details. We have also enlarged the details to make it easier for readers to compare. As shown in Figure 3 (a), our method demonstrates high-quality audio-visual alignment and acoustic quality relative to state-of-the-art dubbing baseline. In the corresponding silent video frames (see green arrows), our method can generate the same sound pauses as GT, which illustrates the effectiveness of dual contrastive aligning. As shown in Figure 3 (b), we visualize the mel-spectrogram generation effect of Acoustics Flow Matching Guidance at different scales. As the scale increases, the originally blurry and artifact-filled spectrum gradually becomes clearer. The qualitative analysis shows that our model can generate high-quality audio-visual alignment, high-fidelity acoustic quality and speech.
5. Conclusion
In this paper, we propose an LLM-based dubbing architecture, which incorporates speech large language model for semantic-aware learning and voice enhanced flow matching for high-quality acoustic modeling. By LLM-based semantic-aware learning, the model absorbs the phoneme level semantic knowledge with in-contextual information, while maintaining the lip-sync with silent video by dual contrastive aligning. Besides, the proposed Flow-based Voice Enhancing ensures the acoustic clarity and speaker identity by LLM-based Acoustics Flow Matching Guidance and style flow matching prediction. Our proposed model sets new state-of-the-art on both Chem and GRID benchmarks.
Appendix A Theoretic Details of Flow Matching
Given the mel-spectrograms data space with data point $M$, where $M\sim q(M)$ and $q(M)$ is an unknown data distribution of mel-spectrograms, a possible approach to sample $M$ from $q(M)$ is from the probability density path defined as $p_{t}(x)$ where $t\in[0,1]$, $p_{0}(x)=\mathcal{N}(x;\boldsymbol{0},\boldsymbol{I})$ and $p_{1}(x)\approx q(x)$. To estimate the probability density path, Continuous Normalizing Flow (Chen et al., 2018) defines a vector field $v_{t}$, which gives a flow $\phi_{t}(x)$ through an Ordinary Differential Equation (ODE):
(13) $\frac{d}{dt}\phi_{t}(x)=v_{t}(\phi_{t}(x));\quad\phi_{0}(x)=x.$
Chen et al. (2018) shows that predicting the vector flow $v_{t}$ through a neural network $\theta$ can be used to transform a simple Gaussian distribution to a more complicated one, such as $q(x)$, which generates the probability density path. Flow matching is designed to predict such a probability density path. Give a target probability density path $p_{t}$ with its corresponding known vector field $u_{t}(x)$, the training objective of flow matching is:
(14) $\mathcal{L}_{FM}(\theta)=\mathbb{E}_{t,p_{t}(x)}||v_{t}(x,\theta)-u_{t}(x)||^{2}$
However, the total probability density path $p_{t}$ is unknown and we can only use some samples from $q(x)$ to estimate the probability density path, which is Conditional Flow Matching (CFM). The training objective of CFM is:
(15) $\mathcal{L}_{CFM}(\theta)=\mathbb{E}_{t,M\sim q(M),p_{t}(x|M)}||v_{t}(x,\theta)-u_{t}(x|M)||^{2},$
We can efficiently estimate the probability density path $p_{t}(x)$ by sampling from $q(M),p_{t}(x|M)$ and calculate $u_{t}(x|M)$. Here we use optimal-transport conditional flow matching (OT-CFM) to train our model, which is a simple version of CFM with simple flow $\phi_{t}(x)=(1-(1-\sigma_{\min})t)x_{0}+tM$, which satisfies $x_{0}\sim\mathcal{N}(x;\boldsymbol{0},\boldsymbol{I})$ and $\phi_{1}(x)\sim q(x)$. Its gradient vector field is $u_{t}(\phi_{t}(x)|M)=M-(1-\sigma_{min})x_{0}$, enabling fast training and inference for its linear and time-invariant properties. Given conditional mean $\mu$, the training objective of OT-CFM can be formulated as:
(16) $\mathcal{L_{\theta}}=\mathbb{E}_{t,q(M),p_{t}(x|\mu,M)}||v_{t}(\phi_{t}(x)|\mu,\theta)-u_{t}(\phi_{t}(x)|M)||^{2},$
where $v_{t}(\phi_{t}(x)|\mu,\theta)$ is the predicted gradient vector field of $\phi_{t}(x)$ according to the acoustics prior information $\mu$. Then, we can solve the ODE $d\phi_{t}(x)=v_{t}(\phi_{t}(x)|\mu,\theta)dt$ from $t=0$ to $t=1$ to generate the target mel-spectrogram $\hat{M}$ from noise $x_{0}$.
Table 8. Subjective evaluation on GRID benchmark.
Dataset GRID
Methods MOS-N $\uparrow$ MOS-S $\uparrow$ CMOS $\uparrow$
GT 4.69$\pm$0.07 - +0.10
V2C-Net (Chen et al., 2022) 3.62$\pm$0.06 3.67$\pm$0.11 -0.35
HPMDubbing (Cong et al., 2023) 3.77$\pm$0.20 3.74$\pm$0.13 -0.26
StyleDubber (Cong et al., 2024b) 4.02$\pm$0.11 4.06$\pm$0.05 -0.19
Speaker2Dubber (Zhang et al., 2024) 4.10$\pm$0.09 4.05$\pm$0.11 -0.18
Produbber (Zhang et al., 2025a) 4.12$\pm$0.07 4.07$\pm$0.10 -0.13
Our method 4.15$\pm$0.06 4.10$\pm$0.07 0.00
Appendix B LSE or MCD-DTW-SL for measuring lip-sync?
MCD-DTW-SL cannot truly measure audiovisual synchronization because the coefficients of MCD-DTW-SL are based on the global time rather than the fact to reflect the alignment related to lip movement. Below is the formula of MCD-DTW-SL:
[…]
(18) $n=\frac{T_{mel}}{T_{v}}=\frac{sr/hs}{FPS}\in\mathbb{N}^{+},$
where $FPS$ denotes the Frames per Second of the video, $sr$ denotes the sampling rate of the audio, and $hs$ denotes hop size when transforming the raw waveform into mel-spectrograms. In this case, the audio length is known in advance, MCD-DTW-SL is meaningless for determining alignment. Thus, we encourage evaluating the audio-visual synchronization in movie dubbing by using metrics LSE-D and LSE-C, which are widely adopted for quantitative evaluation of lip-syncing performance in the wild (Yaman et al., 2024; Ki and Min, 2023; Yang et al., 2024a; Tan et al., 2024; Park et al., 2024). The LSE-D measures the distance between the audio and visual representations with lower scores suggesting better audio-visual sync. LSE-C is the confidence score, and the higher value implies a stronger correlation between video and speech (Lei et al., 2024; Liu et al., 2024; Sahipjohn et al., 2024).
Appendix C Subjective Evaluation on GRID Benchmark
Since we only provide objective evaluation in the main paper, we provide the subjective evaluation by human, following previous dubbing works (Zhang et al., 2024, 2025a). Please note that the UTMOS and DNSMOS in the paper are given by the model prediction to ensure fairness, rather than human subjective evaluation.
MOS-N & MOS-S. MOS-Naturalness (MOS-N) and MOS-Similarity (MOS-S) are mean option scores reported with a 95% confidence interval based on ratings from 20 native English speakers using a scale from 1 to 5. Each participant is required to listen to 30 randomly selected generated dubbing and rate the dubbing according to the speech naturalness and voice similarity following (Chen et al., 2022).
CMOS. Comparative mean option score (CMOS) asks participants to compare the dubbing generated by two models using same input and rate them on a scale from -5 to 5 based on criterion of matched degree between generated dubbing with video (Zhang et al., 2025a).
Under the same experimental setting, we found that the proposed FlowDubber is also subjectively superior to the previous methods, especially in speech quality (MOS-N), indicating the best overall dubbing quality.